Calculate the consumer surplus for the demand function P = 25/Q+2 when the market price P = 5.


1
Expert's answer
2021-07-11T17:39:01-0400

We use the equilibrium price provided to solve for Q.

Therefore,


"5=\frac{25}{Q+2}"


"Q=3" when we simplify the equation.


In order to calculate the Consumer Surplus, we need to integrate the demand function at Q=3 and Q=0 range. We would get 25−Q2dQ−(5)(3)


"25Q-\frac{Q^3}{3}-15"


we substitute Q with 3 from the equation.


"(25\times3)-\frac{3^3}{3}-15"


51


Therefore Consumer Surplus will be $51